# Gravitational Waves from Massive Black Hole Mergers in ASTRID: Predictions for LISA

### v1

@@ 3.1 MBH population in ASTRID
In Figure 3, we present an illustration of LISA sources in a slice of $100\times 60\times 10\ \mathrm{Mpc}^{3}\,h^{-3}$ extracted from ASTRID at $z=0.8$. The upper panel illustrates the gas density field color-coded by the temperature, with warmer regions corresponding to higher temperatures. Yellow circles mark the locations of the mergers occurring within a 670 Myr period centered on $z=0.8$. We select five merging systems as examples, which cover the whole mass range ($10^{5}\sim 10^{8}\ \mathrm{M}_{\odot}$) for the LISA sources, and mark their position with arrows. The middle panels show the stellar density field of the host halo for the five systems, color-coded by stellar age, where blue regions represent young stars. The red crosses show the position of the primary MBH (or the remnant MBH), and the green crosses show the secondary MBH. The bottom panels show their host galaxy, with the RGB channels representing the flux in the rest-frame $grz$ color bands.
@@
In Figure 5 we present the merging orbit eccentricity $e$ distributions, which are estimated based on the method described in Section 2.2. The left panel shows the probability density function (PDF) for $e$ at different redshifts. Across the whole redshift range, $e$ peaks at high values: $e\gtrsim 0.8$, and it evolves to a smaller value at lower redshift: from $0.88$ at $z>3$ to $0.80$ at $z<0.5$, indicating a larger fraction of binaries at low redshift merge with circular orbits. A potential explanation for this evolution is that binaries with eccentric orbits have shorter hardening timescales (Chen et al., 2022b) and they will merge shortly after the binaries are formed, while many circular binaries keep evolving towards lower redshift. To give a better description of the $e$ distribution, we fit the PDF shown in Fig. 5 using a $\beta$ distribution:
@@
The normalization factor is $\mathcal{N}=1/\mathrm{B}(\alpha,\beta)$, where $\mathrm{B}\left(\alpha,\beta\right)$ is the $\beta$ function. The two shape parameters $\alpha$ and $\beta$ depend linearly on redshift $z$: $\alpha=1.08\,(1+z)+2.08$, $\beta=0.07\,(1+z)+1.56$. The fitted distributions are plotted in Figure 5 with dotted curves. The right panel shows the PDF within different total MBH masses. Although the evolution of $e$ over $M_{\mathrm{BH}}$ is less obvious compared to that over $z$, we note that the most massive mergers $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ typically have higher $e$.
@@ 3.2 LISA Predictions
In this subsection, we investigate the MBH mergers in the Circ/Ecc run that are detectable by LISA (i.e., with SNR $>10$), and make predictions for the LISA detection rate.
Figure 8: The LISA detection rate estimated using 10,000 realizations of MBH population. $Upper:$ the number of observed binaries (with SNR $>10$) per year per unit redshift (left) / log $M_{\mathrm{BH}}$ (right). In each panel, the yellow/blue solid curves represent the results from Circ/Ecc run, respectively. The blue dashed curve shows the predicted inspiral sources from the Ecc runs. We do not plot the inspiral sources for the Circ since it has a negligible number of inspiral sources (less than $10^{-3}$/yr). In the upper left frame, the vertical dotted line represents the peak of redshift distribution: $z_{\mathrm{peak}}=1.9$ for Circ (yellow) and $z_{\mathrm{peak}}=0.80$ (blue) for Ecc run. $Bottom:$ the fraction between the number of detectable sources in Circ and Ecc. We mark the level of $N_{\mathrm{Ecc}}=N_{\mathrm{Circ}}$ with the black dashed line.
@@
We take the waveforms calculated in Section 2.3, and compute the strain and frequency distributions for the MBH mergers in ASTRID. The results are shown in Fig. 6, where the left panel is for the Circ run and the right panel for the Ecc run. For each MBH merger, we find the point in its time evolution and the harmonic mode that makes the greatest contribution to the SNR in the 4 years before coalescence. This is used to position the source in Fig. 6. For the Circ run, the harmonic mode is always $n=2$. We color the distribution based on the redshift and the number of events, as shown by the right colorbar: the high (low) redshifts are represented by blue (red) colors, and the darkness corresponds to the regions with very few events. The redshift is averaged over the merger population within each pixel. The cut around $f\sim 10^{-2.5}$ Hz corresponds to the LISA sensitivity peak, which arises from our choice of plotting the frequency that contributes the most to the SNR. As we can see, the strain-frequency distributions for the Circ and Ecc runs are generally similar, except that the scatter in the strain range for a specific redshift is slightly larger for Ecc run.
In Fig. 7, we demonstrate the parameter space of MBH mass and redshift explored by LISA. The left panel shows the two-dimensional distribution of redshift and total merger mass $M_{\mathrm{tot}}$, and the right panel shows how the mass of the primary MBH $M_{1}$ and secondary MBH $M_{2}$ are distributed. The underlying blue color gradient represents the differential number of the detected mergers during the LISA 4-year observation for Ecc run: $dN/(d\log M\,dz\,dt)\times\mathrm{4yr}$ for the left panel and $dN/(d\log M_{1}\,d\log M_{2}\,dt)\times\mathrm{4yr}$ for the right panel. The black and red contours correspond to the detected MBH merger population for Ecc run and Circ run, respectively. We can see that the detected eccentric and circular binaries largely overlap, and most of the LISA sources are low-mass mergers with $M_{\mathrm{tot}}<10^{7}$ $\mathrm{M}_{\odot}$ at the local Universe ($z<2$). An important difference between Ecc and Circ observed in Fig. 7 is that including eccentricity enables LISA to explore sources with higher mass. Around $z\sim 0$, the most massive mergers within detectable circular binaries have mass $\sim 10^{8}$ $\mathrm{M}_{\odot}$, while for eccentricity sources, the high-mass limit increases to $M_{\mathrm{tot}}\sim 10^{9}$ $\mathrm{M}_{\odot}$. This is also supported by the $M_{2}-M_{1}$ distribution plot: in Ecc run, the covered mass spectra is increased by one order of magnitude compared to the Circ run.
@@
An important message we can get from Fig. 8 is that incorporating orbital eccentricity boosts the LISA detectability. Such an enhancement comes almost fully from the inspiral sources: the number of detected merger sources is almost the same in the Ecc run (5.7/yr) and in the Circ run (5.6/yr). In the Ecc run, at low redshift ($z<0.5$) or at the high mass end $M>10^{7.5}$, the detected inspiral sources outnumber the merger sources by almost an order of magnitude. For the inspiral sources, as the frequency evolution during the early stages is typically slow (Sesana, 2010), the covered frequency range for an individual harmonic mode is very narrow, producing a smaller SNR for the inspiral sources in the Circ run. On the other hand, in the Ecc run, the harmonic modes higher than $n>2$ not only result in a wider frequency range but also extend the signal to the band where LISA has high sensitivity (see the right panel of Fig. 2). This makes the inspiral sources more likely to be detected when the eccentricity is taken into account.
@@
A caveat is that at $z>2$ or $M_{\mathrm{BH}}<10^{6}$ $\mathrm{M}_{\odot}$, the number of detectable inspiral sources drops quickly. This is because at high redshift, or the low mass end, the strain is too low to produce a detectable signal even if the high harmonic mode is close to the LISA high-sensitivity band. Moreover, for the eccentricity binaries, their peak strain is lower than their circular counterpart, which can be seen from Fig. 2 (middle and right panels). This explains why at high redshift ($z>6$), Circ run predicts a slightly larger detection rate than the Ecc run. The mass peak for the inspiral sources at $M_{\mathrm{tot}}\sim 10^{6}$ $\mathrm{M}_{\odot}$ shifts the entire distribution of detectable sources in the Ecc run to slightly higher mass. In the Circ run, where almost all the sources are merger sources, the mass peak is around the black hole seed mass $M_{\mathrm{tot}}\sim 10^{5}$ $\mathrm{M}_{\odot}$.
@@
The minor increase in the number of detected merger sources in the Ecc run (from $5.6$/yr in Circ run to $5.7$/yr) is from massive mergers. With the circular orbit assumption, no mergers with $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ can be detected, while in the Ecc run, there are still some sources with $M_{\mathrm{BH}}\gtrsim 10^{9}$ $\mathrm{M}_{\odot}$ observed through coalescence in the 4-year observation. This corresponds to the gap between the blue solid curve and the blue dashed curve. Binaries with such a large mass typically merge at low frequency before they are observed in the LISA band. When we include the eccentricity, the higher harmonic modes increase the emission at higher frequency, which is closer to the LISA high-sensitivity band, and this contributes significantly to the total SNR. This is the case plotted in the right panel of Fig. 2.
@@
In Fig. 9, we present the detection rate as a function of the MBH mass ratio $q=M_{2}/M_{1}$. In the upper panel, besides all the detected sources in Ecc run (blue solid curve), Circ run (yellow solid curve), and the inspiral sources in the Ecc run (blue dashed curve), we also plot the total merger rate in ASTRID (black dotted curve). Most of the LISA sources have a low mass ratio: $q<0.1$. This is because there are more low-$q$ mergers in ASTRID but not because LISA is more sensitive to them. Actually, LISA can observe almost all the mergers with $q\geq 0.1$, which is shown by the overlap between the yellow curve and black curve, while a large fraction of low-$q$ mergers is not detectable. Note that the mass ratio has a dependence on the merger mass $M_{\mathrm{BH}}$: low-$q$ mergers are dominated by massive mergers, since the secondary MBH mass $M_{2}$ is still massive than the adopted black hole seed mass $10^{4.5}$ $\mathrm{M}_{\odot}$. This dependence explains why the detection rate for the Circ run drops at low-$q$ end, and the enhancement of the detection rate in Ecc run at $q<0.1$ is consistent with the right panel of Fig. 8.
@@
Figure 10: The galaxies hosting detectable sources at $z=0.8$. The contour plots in the leftmost column compare the properties of host galaxies of detectable LISA sources which merger within $0.90.8<z<$ 0.9 (red contours) and the galaxies hosting no mergers (grey contours). Each contour set plots the $1\sigma/2\sigma/3\sigma$ regions. Left Top: Merger remnant MBH mass versus galaxy stellar mass. Left Middle: Star formation rate versus galaxy stellar mass. Note that we remove the galaxies with zero star formation rate to plot this distribution. Left Bottom: Merger remnant MBH mass versus the stellar mass ratio between the host galaxy and the FoF halo ($M_{\rm gal}/M_{\rm FOF}$). In the top and bottom panels, the non-merging galaxy population uses the central MBH mass for the y-axis as a comparison. We randomly select a sample consisting of 6 detectable sources, which are marked by the diamonds in the contours plots. In the middle and right columns, we plot the stellar density field for the FoF halo for the selected mergers, the location of the remnant MBH is marked by the arrows with corresponding colors. In each frame, the circle represents the virial radius of the halo.
@@ 4 Host Galaxy and Environment
Figure 12: 1-D histograms presenting different properties of merging black holes and their subhalo environments. Blue and red correspond to redshifts z = 1.5 and z = 0.8, respectively. The smooth traces correspond to the distributions of merger events with SNR $>$ 1000. Left: Ratio of total merger mass ($M_{1}+M_{2}$) to black hole mass in the corresponding subhalo. Middle: Black hole Eddington ratio. Right: Black hole Bolometric luminosity.
An important advantage of using MBH populations from cosmological simulations like ASTRID over those generated by the semi-analytical models (SAMs) is that we can self-consistently associate the MBH merger with their host galaxy. In this section, we focus on the merging populations at two redshifts, $z=1.5$ and $z=0.8$ to investigate the correlation between the mergers and their local environments. We use the SNR calculated in the Circ run to determine the detectable merger population in this section.
@@ 4.1 Host galaxies of detected sources
Identifying host galaxies or AGN activities of gravitational wave sources is key to the multi-messenger study of MBHs and has the potential for constraining cosmology. Future detectors like LISA will be able to localize the GW sources to the square-degree level, much better than the current ground-based detectors. However, there are still up to a few hundred candidate galaxies in the LISA source field (Lops et al., 2023). Being able to distinguish the merger hosts from single-MBH galaxies greatly facilitates the EM-followup of LISA events. Recently, Bardati et al. (2024a) and Bardati et al. (2024b) investigated the morphological signatures of MBH merger hosts using cosmological simulation Romulus and found persistent morphological changes in high-mass major mergers. Izquierdo-Villalba et al. (2023) investigated the signatures of LISA source hosts using SAM and found it difficult to identify the merger hosts solely from the galaxy properties. In this section, we look for distinctive features of MBH merger hosts from the wide range of galaxy population from ASTRID.
@@
In Fig. 10 and 11 we present the properties of the galaxies hosting LISA sources, and compare them to the whole galaxy population in ASTRID at $z=0.8$ and $z=1.5$, respectively. In the left column of both figures, the red contours represent the galaxy hosting the MBH mergers detectable by the LISA, i.e., SNR $>10$, whose boundaries are highlighted by the dashed lines; and the contours with a grey color scheme correspond to the galaxy population not hosting mergers at that redshift. For each contour set, we plot the $1\sigma/2\sigma/3\sigma$ regions. From top to bottom, we show the two-dimensional distribution of the MBH mass versus galaxy stellar mass ($M_{\mathrm{BH}}-M_{\mathrm{gal}}$, top), galaxy star-formation rate versus galaxy stellar mass ($\mathrm{SFR}-M_{\mathrm{gal}}$, middle), and MBH mass versus the stellar mass ratio between the host galaxy and host Friends-of-Friends (FoF) halo ($M_{\mathrm{BH}}-M_{\mathrm{gal}}/M_{\mathrm{FoF}}$, bottom). For the galaxies hosting detectable sources, we plot the mass of the remnant MBH as $M_{\mathrm{BH}}$; and for the non-merger galaxies, we plot the mass of the central MBH. At both redshifts, we randomly select a sample consisting of six detectable sources, whose positions among the two-dimensional contours are marked by diamonds. In the middle and right columns, we plot the stellar density field for the FoF halo for the selected mergers. The location of the remnant MBH is marked by symbols with colors corresponding to those used for the same object in the left panels. In each frame, the circle represents the virial radius of the halo.
[…]
$\log(\mathrm{SFR})\;=\;\alpha\,\log\bigl{(}M_{\mathrm{gal}}\bigr{)}\;+\;\beta.$ (10)
For $z=0.8$ we find that the slope $\alpha=0.88\pm 0.10$ and the intercept $\beta=-8.07\pm 0.90$. The scatter about this relationship is indeed quite small, $\sigma_{\mathrm{rms}}\approx 0.17\,\mathrm{dex}$. The fit values for $z=1.5$ are almost the same, and consistent within the errors: $\alpha=0.90\pm 0.09$, $\beta=-8.04\pm 0.83$, and $\sigma_{\mathrm{rms}}\approx 0.14\,\mathrm{dex}$.
Some merger hosts have SFR of 10-100 $\mathrm{M}_{\odot}$/yr, and if we examine two of these (the orange and green symbols in Fig. 10 we can see that while both lie at the centers of massive galaxies, one (orange) is in an obviously disturbed ongoing galaxy merger, whereas the green symbol is in a much more uniform and symmetric galaxy. Our merger galaxy masses are similar to those predicted by recent work using SAMs (e.g. Izquierdo-Villalba et al., 2023), but we found a higher contrast in the SFR between merging and single MBH hosts in our simulation.
@@
Looking at the bottom left panels of Figure 10 and 11 that merger hosts are more likely to be the most massive galaxy in a halo. For example, 68% of all merger hosts have a mass greater than $\sim 20\%$ of that of the FOF halo they reside in, but 68% of non-mergers lie in satellite galaxies with mass less than 0.5% of their FOF halo. Having said this, however, a significant fraction of the merger hosts ($\sim$ 20-30%) do lie in satellite galaxies in the outskirts of massive halos. For example, among the six examples we have selected and plotted, only half of them are located in the centers of their parent halos, and one of them at each redshift is even hosted by galaxies outside the virial radius (the indigo diamond in Figure 10 and pink diamond in Figure 11). The high probability of merger hosts in isolated halos compared with single MBH hosts may facilitate the EM identification of the LISA sources among the $>100$ candidate galaxies in the LISA source field (e.g. Lops et al., 2023).

### v2

@@ 3.1 MBH population in ASTRID
In Figure 4, we present an illustration of LISA sources in a slice of $100\times 60\times 10\ \mathrm{Mpc}^{3}\,h^{-3}$ extracted from ASTRID at $z=0.8$. The upper panel illustrates the gas density field color-coded by the temperature, with warmer regions corresponding to higher temperatures. [^3] The visualization was created using Blender software; see the companion video at https://astrid.psc.edu/gallery/ Yellow circles mark the locations of the mergers occurring within a 670 Myr period centered on $z=0.8$. We select five merging systems as examples, which cover the whole mass range ($10^{5}\sim 10^{8}\ \mathrm{M}_{\odot}$) for the LISA sources, and mark their position with arrows. The middle panels show the stellar density field of the host halo for the five systems, color-coded by stellar age, where blue regions represent young stars. The red crosses show the position of the primary MBH (or the remnant MBH), and the green crosses show the secondary MBH. The bottom panels show their host galaxy, with the RGB channels representing the flux in the rest-frame $grz$ color bands.
@@
In Figure 5 we present the merging orbit eccentricity $e_{0}$ distributions, which are estimated based on the method described in Section 2.2. The left panel shows the probability density function (PDF) for $e_{0}$ at different redshifts. Across the whole redshift range, $e_{0}$ peaks at high values: $e_{0}\gtrsim 0.8$, and it evolves to a smaller value at lower redshift: from $0.88$ at $z>3$ to $0.80$ at $z<0.5$, indicating a larger fraction of binaries at low redshift merge with circular orbits. A potential explanation for this evolution is that binaries with eccentric orbits have shorter hardening timescales (Chen et al., 2022b) and they will merge shortly after the binaries are formed, while many circular binaries keep evolving towards lower redshift. To give a better description of the $e_{0}$ distribution, we fit the PDF shown in Fig. 5 using a $\beta$ distribution:
@@
The normalization factor is $\mathcal{N}=1/\mathrm{B}(\alpha,\beta)$, where $\mathrm{B}\left(\alpha,\beta\right)$ is the $\beta$ function. The two shape parameters $\alpha$ and $\beta$ depend linearly on redshift $z$: $\alpha=1.08\,(1+z)+2.08$, $\beta=0.07\,(1+z)+1.56$. The fitted distributions are plotted in Figure 5 with dotted curves. The right panel shows the PDF within different total MBH masses. Although the evolution of $e_{0}$ over $M_{\mathrm{BH}}$ is less obvious compared to that over $z$, we note that the most massive mergers $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ typically have higher $e_{0}$.
@@ 3.2 LISA Predictions
In this subsection, we investigate the MBH mergers in the Circ/Ecc that are detectable by LISA (i.e., with SNR $>10$), and make predictions for the LISA detection rate.
Figure 8: The LISA detection rate estimated using 10,000 realizations of MBH population. $Upper:$ the number of observed binaries (with SNR $>10$) per year per unit redshift (left) / log $M_{\mathrm{BH}}$ (right). In each panel, the yellow/blue solid curves represent the results from Circ/Ecc, respectively. The blue dashed curve shows the predicted inspiral sources from the Ecc. We do not plot the inspiral sources for the Circ since it has a negligible number of inspiral sources (less than $10^{-3}$/yr). In the upper left frame, the vertical dotted line represents the peak of redshift distribution: $z_{\mathrm{peak}}=1.9$ for Circ (yellow) and $z_{\mathrm{peak}}=0.80$ (blue) for Ecc. $Bottom:$ the fraction between the number of detectable sources in Circ and Ecc. We mark the level of $N_{\mathrm{Ecc}}=N_{\mathrm{Circ}}$ with the black dashed line.
@@
We take the waveforms calculated in Section 2.3, and compute the strain and frequency distributions for the MBH mergers in ASTRID. The results are shown in Fig. 6, where the left panel is for the Circ and the right panel for the Ecc. For each MBH merger, we find the point in its time evolution and the harmonic mode that makes the greatest contribution to the SNR in the 4 years before coalescence. This is used to position the source in Fig. 6. For the Circ, the harmonic mode is always $n=2$. We color the distribution based on the redshift and the number of events, as shown by the right colorbar: the high (low) redshifts are represented by blue (red) colors, and the darkness corresponds to the regions with very few events. The redshift is averaged over the merger population within each pixel. The cut around $f\sim 10^{-2.5}$ Hz corresponds to the LISA sensitivity peak, which arises from our choice of plotting the frequency that contributes the most to the SNR. As we can see, the strain-frequency distributions for the Circ and Ecc are generally similar, except that the scatter in the strain range for a specific redshift is slightly larger for Ecc.
In Fig. 7, we demonstrate the parameter space of MBH mass and redshift explored by LISA. The left panel shows the two-dimensional distribution of redshift and total merger mass $M_{\mathrm{tot}}$, and the right panel shows how the mass of the primary MBH $M_{1}$ and secondary MBH $M_{2}$ are distributed. The underlying blue color gradient represents the differential number of the detected mergers during the LISA 4-year observation for Ecc: $dN/(d\log M\,dz\,dt)\times\mathrm{4yr}$ for the left panel and $dN/(d\log M_{1}\,d\log M_{2}\,dt)\times\mathrm{4yr}$ for the right panel. The black and red contours correspond to the detected MBH merger population for Ecc and Circ, respectively. We can see that the detected eccentric and circular binaries largely overlap, and most of the LISA sources are low-mass mergers with $M_{\mathrm{tot}}<10^{7}$ $\mathrm{M}_{\odot}$ at the local Universe ($z<2$). An important difference between Ecc and Circ observed in Fig. 7 is that including eccentricity enables LISA to explore sources with higher mass. Around $z\sim 0$, the most massive mergers within detectable circular binaries have mass $\sim 10^{8}$ $\mathrm{M}_{\odot}$, while for eccentricity sources, the high-mass limit increases to $M_{\mathrm{tot}}\sim 10^{9}$ $\mathrm{M}_{\odot}$. This is also supported by the $M_{2}-M_{1}$ distribution plot: in Ecc, the covered mass spectra is increased by one order of magnitude compared to the Circ. Chen et al. (2025) found the bulk of the GW background detected by PTAs originates from MBH binaries with masses $M_{\mathrm{tot}}=10^{9}\sim 3\times 10^{9}$ $\mathrm{M}_{\odot}$ at low-frequency $f=4$ nHz, and from binaries within $M_{\rm tot}=10^{7.5}\sim 10^{9.3}$ $\mathrm{M}_{\odot}$at high-frequency $f=20$ nHz. Our results in Fig. 7 implies that some LISA events may also be PTA sources.
@@
An important message we can get from Fig. 8 is that incorporating orbital eccentricity boosts the LISA detectability. Such an enhancement comes almost fully from the inspiral sources: the number of detected merger sources is almost the same in the Ecc (5.7/yr) and in the Circ (5.6/yr). In the Ecc, at low redshift ($z<0.5$) or at the high mass end $M>10^{7.5}$, the detected inspiral sources outnumber the merger sources by almost an order of magnitude. For the inspiral sources, as the frequency evolution during the early stages is typically slow (Sesana, 2010), the covered frequency range for an individual harmonic mode is very narrow, producing a smaller SNR for the inspiral sources in the Circ. On the other hand, in the Ecc, the harmonic modes higher than $n>2$ not only result in a wider frequency range but also extend the signal to the band where LISA has high sensitivity (see the right panel of Fig. 2). This makes the inspiral sources more likely to be detected when the eccentricity is taken into account.
@@
A caveat is that at $z>2$ or $M_{\mathrm{BH}}<10^{6}$ $\mathrm{M}_{\odot}$, the number of detectable inspiral sources drops quickly. This is because at high redshift, or the low mass end, the strain is too low to produce a detectable signal even if the high harmonic mode is close to the LISA high-sensitivity band. Moreover, for the eccentricity binaries, their peak strain is lower than their circular counterpart, which can be seen from Fig. 2 (middle and right panels). This explains why at high redshift ($z>6$), Circ predicts a slightly larger detection rate than the Ecc. The mass peak for the inspiral sources at $M_{\mathrm{tot}}\sim 10^{6}$ $\mathrm{M}_{\odot}$ shifts the entire distribution of detectable sources in the Ecc to slightly higher mass. In the Circ, where almost all the sources are merger sources, the mass peak is around the black hole seed mass $M_{\mathrm{tot}}\sim 10^{5}$ $\mathrm{M}_{\odot}$.
@@
The minor increase in the number of detected merger sources in the Ecc (from $5.6$/yr in Circ to $5.7$/yr) is from massive mergers. With the circular orbit assumption, no mergers with $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ can be detected, while in the Ecc, there are still some sources with $M_{\mathrm{BH}}\gtrsim 10^{9}$ $\mathrm{M}_{\odot}$ observed through coalescence in the 4-year observation. This corresponds to the gap between the blue solid curve and the blue dashed curve. Binaries with such a large mass typically merge at low frequency before they are observed in the LISA band. When we include the eccentricity, the higher harmonic modes increase the emission at higher frequencies. This is the case plotted in the right panel of Fig. 2.
@@
In Fig. 9, we present the detection rate as a function of the MBH mass ratio $q=M_{2}/M_{1}$. In the upper panel, besides all the detected sources in Ecc (blue solid curve), Circ (yellow solid curve), and the inspiral sources in the Ecc (blue dashed curve), we also plot the total merger rate in ASTRID (black dotted curve). Most of the LISA sources have a low mass ratio: $q<0.1$. This is because there are more low-$q$ mergers in ASTRID but not because LISA is more sensitive to them. Actually, LISA can observe almost all the mergers with $q\geq 0.1$, which is shown by the overlap between the yellow curve and black curve, while a large fraction of low-$q$ mergers is not detectable. Note that the mass ratio has a dependence on the merger mass $M_{\mathrm{BH}}$: low-$q$ mergers are dominated by massive mergers, since the secondary MBH mass $M_{2}$ is still massive than the adopted black hole seed mass $10^{4.5}$ $\mathrm{M}_{\odot}$. This dependence explains why the detection rate for the Circ drops at low-$q$ end, and the enhancement of the detection rate in Ecc at $q<0.1$ is consistent with the right panel of Fig. 8.
@@
Figure 10: The galaxies hosting detectable sources at $z=0.8$. The top contour plots compare the properties of host galaxies of detectable LISA sources which merger within $0.90.8<z<$ 0.9 (red contours) and the galaxies hosting no mergers (grey contours). Each contour set plots the $1\sigma/2\sigma/3\sigma$ regions, enclosing 68%/95%/99.7% of the population. Top Left: The MBH mass versus galaxy stellar mass ($M_{\mathrm{BH}}-M_{\mathrm{gal}}$). Top Middle: The galaxy star-formation rate versus galaxy stellar mass ($\mathrm{SFR}-M_{\mathrm{gal}}$). Note that we remove the galaxies with zero star formation rate to plot this distribution. Top Right: The MBH mass versus the stellar mass ratio between the host galaxy and host Friends-of-Friends (FoF) halo ($M_{\mathrm{BH}}-M_{\mathrm{gal}}/M_{\mathrm{FoF}}$). For the galaxies hosting detectable sources, we plot the mass of the remnant MBH as $M_{\mathrm{BH}}$; and for the non-merger galaxies, we plot the mass of the central MBH. We randomly select a sample consisting of 6 detectable sources, which are marked by the diamonds in the contour plots. In the second and third rows, we plot the stellar density field for the FoF halo for the selected mergers. The location of the remnant MBH is marked by the arrows with corresponding colors. In each frame, the circle represents the virial radius of the halo.
@@ 4 Host Galaxy and Environment
An important advantage of using MBH populations from cosmological simulations like ASTRID over those generated by the semi-analytical models (SAMs) is that we can self-consistently associate the MBH merger with their host galaxy. In this section, we focus on the merging populations at two redshifts: $z=1.5$ and $z=0.8$ to investigate the correlation between the mergers and their local environments.
@@ 4.1 Host galaxies of detected sources
Identifying host galaxies or AGN activities of gravitational wave sources is key to the multi-messenger study of MBHs and has the potential for constraining cosmology. Future detectors like LISA will be able to localize the GW sources to the square-degree level, much better than the current ground-based detectors. However, there are still up to a few hundred candidate galaxies in the LISA source field (Lops et al., 2023). Being able to distinguish the merger hosts greatly facilitates the EM-followup of LISA events. Recently, Bardati et al. (2024a) and Bardati et al. (2024b) investigated the morphological signatures of MBH merger hosts using cosmological simulation Romulus (Tremmel et al., 2017) and found persistent morphological changes in high-mass major mergers. Izquierdo-Villalba et al. (2023) investigated the signatures of LISA source hosts using SAM and found it difficult to identify the merger hosts solely from the galaxy properties. In this section, we look for distinctive features of MBH merger hosts from the wide range of galaxy population from ASTRID.
@@
In Fig. 10 and 11 we present the properties of the galaxies hosting LISA sources, and compare them to the galaxy population hosting no mergers at $z=0.8$ and $z=1.5$, respectively. In the top row of both figures, the red contours represent the galaxy hosting the MBH mergers detectable by the LISA, i.e., SNR $>10$, whose boundaries are highlighted by the dashed lines; and the contours with a grey color scheme correspond to the galaxy population not hosting mergers at that redshift. For each contour set, we plot the $1\sigma/2\sigma/3\sigma$ regions, enclosing 68%/95%/99.7% of the population. From left to right, we show the two-dimensional distribution of the MBH mass versus galaxy stellar mass ($M_{\mathrm{BH}}-M_{\mathrm{gal}}$, top left), galaxy star-formation rate versus galaxy stellar mass ($\mathrm{SFR}-M_{\mathrm{gal}}$, top middle), and MBH mass versus the stellar mass ratio between the host galaxy and host Friends-of-Friends (FoF) halo ($M_{\mathrm{BH}}-M_{\mathrm{gal}}/M_{\mathrm{FoF}}$, top right). For the galaxies hosting detectable sources, we plot the mass of the remnant MBH as $M_{\mathrm{BH}}$; and for the non-merger galaxies, we plot the mass of the central MBH. At both redshifts, we randomly select a sample consisting of six detectable sources, whose positions among the two-dimensional contours are marked by diamonds. In the middle and bottom rows, we plot the stellar density field for the FoF halo for the selected mergers. The location of the remnant MBH is marked by diamonds with the color used for the same object in the contour panel, and the circle represents the virial radius of the halo.
At both $z=0.8$ and $z=1.5$, the galaxies hosting detectable merging sources lie on the same $M_{\mathrm{BH}}-M_{\mathrm{gal}}$ correlation as the non-merger galaxies, but with differences in how they populate this and the rest of the parameter space. Compared to the non-merger galaxy population, merger hosts have larger galaxy mass on average, with the center of the distribution lying at $10^{9}$ $\mathrm{M}_{\odot}$ at both redshifts plotted. This is two orders of magnitude beyond the mean of the non-merger distribution, the latter of which does depend on the lower mass limit for inclusion into the sample. A significant fraction of the merger hosts are also very massive galaxies with large BHs, as the contour enclosing 99.7% of the population extending to $M_{\mathrm{gal}}\sim 10^{12}$ $\mathrm{M}_{\odot}$ and $M_{\mathrm{BH}}\sim 10^{9}$ $\mathrm{M}_{\odot}$. The contours also extend to lower masses, showing that a wide variety of galaxies can be hosts. For example, many detectable mergers of small MBHs take place in galaxies with masses as small as $10^{7}$ $\mathrm{M}_{\odot}$ and below. Our mass range of the merger hosts is similar to those predicted by recent work using SAMs (e.g. Izquierdo-Villalba et al., 2023).
The ${\rm SFR}-M_{\mathrm{gal}}$ distributions are similar among the host galaxies hosting LISA mergers and those that are not. Note however that LISA mergers host galaxies occupy the more massive end of the population. LISA merger host galaxies also have SFRs ranging from $0.07\sim 5.56$ $\mathrm{M}_{\odot}$/yr, ( almost none have a SFR smaller than $10^{-3}$ $\mathrm{M}_{\odot}$/yr, even for the low-mass end. Some merger host galaxies have a high SFR of 10-1000 $\mathrm{M}_{\odot}$/yr, and if we examine two of these (the orange and green symbols in Fig. 10), we can see that while both lie at the centers of massive galaxies, one (shown with orange symbok) is a disturbed ongoing galaxy merger, whereas the other (green) is in a much more uniform (yet somewhat disturbed) galaxy.
@@
The top right panels of Fig. 10 and 11 show that most of the host galaxies of LISA mergers are the central galaxies within a halo. For example, 68% of all merger hosts have a stellar mass greater than $\sim 10\%$ of that of the FOF halo they reside in. However, a significant fraction of the merger hosts ($\sim$ 20-30%) does lie in the satellite galaxies. For example, among the six examples we have selected and plotted, only half of them are located in the centers of their parent halos, and one of them at each redshift is even hosted by galaxies outside the virial radius (the indigo diamond in Fig. 10 and the pink diamond in Fig. 11). The high probability of merger hosts in isolated systems (i.e., with $M_{\mathrm{gal}}/M_{\rm FOF}$ close to 1) may facilitate the EM identification of the LISA sources among the $>100$ candidate galaxies in the LISA source field (e.g. Lops et al., 2023).
@@
Figure 12: The MBH population in the detectable mergers (with SNR $>10$) in the Ecc. Left: Ratio of total merger mass ($M_{1}+M_{2}$) to the total black hole mass in the host galaxies. Middle: MBH Bolometric luminosity $L_{\mathrm{bol}}$. Right: Black hole Eddington ratio $f_{\mathrm{edd}}$. In each frame, the grey histogram represents the distribution of all the mergers across the redshift and mass range. The solid black(red) curves show the MBHs involving mergers with $M_{\mathrm{tot}}>10^{6}$ $\mathrm{M}_{\odot}$($M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$). The dashed histograms correspond to the mergers occurring at high redshift ($z\geq 2$) with $M_{\mathrm{tot}}>10^{6}$ $\mathrm{M}_{\odot}$ (black dashed) or $M_{\mathrm{tot}}>10^{8}$ $\mathrm{M}_{\odot}$ (red dashed).